Acquirer firm plans to launch a takeover of Target firm. The manager of Acquirer indicates that the deal will increase the free cash flow of the combined business by $13.6m per year forever. The beta of the combined firm is 1.2, market portfolio return is 12% and risk free interest rate is 4%. Firms Involved in the Takeover is shown Firms Involved in the Takeover Acquirer Target Assets ($m) 6000 800 Debt ($m) 2000 80 300 20 Number of shares (m) (i). Calculate the value of synergy of the deal. (6 marks) 10 (ii). Calculate the offer price at which Acquirer shares the total synergy as in (i) equally with Target
Answer #1

Re = 1136 Re = of + (rm-of) xß = 4+ (12-4) X112 = 13.6% Cashflow will increase by 13.6m (il Value of Synergy 1316 = 100 m. Va